The focus was off last night. I’ll interrupt the queue at the beginning of tonight to re-calibrate it.
Monthly Archives: June 2012
New UPS ordered
The new building UPS has been ordered and is en route.
A few servicing issues resolved
Brevin went up last night (6/22) to do a quick servicing on the mirror support compressor, the axis video system, and the oculus camera control computer. In short, all issues are fixed, and the telescope is back to normal operations. Please check images from tonight (6/23) to insure that the telescope is functioning properly. Please reply to this message, or email me personally, if you still find issues. Details on each issue are below.
Camera control machine down
The camera control computer is not responding following a loss in power on July 20. Brevin has been alerted.
RCT experienced a power out at 00:10 local time. Telescope was restarted and homed at 11:00 mountain.
Several images the last few nights are blank. I suspect the dome encoders are off, and have re-homed the encoder. Please send comments if your images are still blank.
A note on JPL horizons and non-sidereal requests
There is a bug in the RCT scripts that occurs when parsing names with white spaces from the JPL Horizons ephemeris service. I’m still working on robust solutions. You may try putting quotes around your object names, like “C/2011 L1”. But I recommend looking up targets on http://ssd.jpl.nasa.gov/horizons.cgi to generate a Object Data Page, and using either the Record Number (indicated by Rec #: xxxxx in the upper left hand corner) or the MB number (in the upper right hand corner) as the object name in the IsPlanet key on the RCT Observing Scheduler.
Known issues with PSF quality
Telescope was adjusted for seasonal drift in focus. Images now show a three-point PSF indicicating that the mirror support compressor is not opperating properly.